### v1

@@ 1 Introduction
Semi-empirical Quantum Mechanical methods, such as the Density Functional Tight Binding (DFTB) method, have been used to compute phonons in few highly symmetric isoreticular MOFs [14]. While DFTB is 2–3 orders of magnitude faster than DFT, its pair-wise parametrization strategy and lack of available metal atom parameters limits its use in screening applications for obtaining phonons in MOFs. The GFN1-xTB method [15] emerged to avoid this pair-wise parametrization problem of DFTB and is capable of handling thousands of atoms. In [16], the GFN1-xTB method has been reported to produce cell parameters within a 5% deviation (relative to experiments) for 75% of the MOFs in the CoRE database [17, 18] showing promise for adsorption applications and obtaining binding energies, but has not been analyzed for vibrational properties in MOFs.
Traditional force fields like UFF [19] and CHARMM [20] are widely used due to their scalability and significant speed up in calculations compared to tight-binding DFT. UFF4MOF [21], which is UFF re-parametrized for MOFs, can be used for rapid structure prediction and screening. However, this transferability comes with limitations in accurately predicting dynamical properties such as phonons. [22, 21]. There are other DFT-derived force fields like the MOF-FF model [23], which predicted the lattice parameters of some well known MOFs like MOF-5, UiO-66 within 0.5% deviation relative to DFT. But its system-specific parametrization can be very complex and time-consuming for large phonon calculations in diverse MOFs. Force fields derived with a focus on vibrational properties for MOFs have been reported like VMOF [24]. While the VMOF model reproduced optimized lattice parameters and phonon density of states in good agreement with DFT, important phonon-derived properties like bulk modulus obtained with VMOF underestimated DFT predictions by more than 50% even for standard MOFs like UiO-66 and MOF-5 [24]. These limitations of such classical force fields are because of the vast combinations of possible frameworks in the MOF chemical space that make their parameter optimization and selection of functional form extremely challenging.
@@ 2.1 Dataset Curation
For sampling these 108 structures, we considered MOFs with pore limiting diameters (PLDs) greater than 3.6 Å, using nitrogen as a probe molecule. MACE [34] descriptors hold a large amount of information regarding the atomic features of MOFs and hence were used to sample 100 diverse MOFs. These MOFs were selected to span over all the 7 crystal symmetry systems and diverse bonding interactions. Therefore, as shown in Figure 1 (c, d), this dataset consists of 127 representative MOFs spread over a wide range of 24 elements in the inorganic clusters and organic ligands. Figure 1 (a, b) shows the MACE descriptor space diversity for the curated 127 MOFs (refer the SI for the main features of the sampled MOFs).
@@ 2.2 Phonon Workflow
The Harmonic Approximation (HA) for computing phonons fails to capture essential physical properties for MOFs such as thermal expansion, phase transition and elastic moduli [42]. On the other hand, the Quasi-Harmonic approximation (QHA) is an extension of the HA that aims to capture anharmonicity by introducing the volume dependence of frequencies due to temperature and pressure effects. In this work, we use the Phonopy package version 2.29.0 to study thermal and mechanical properties of MOFs by studying the lattice dynamics under the QHA on a $11\times 11\times 1$ 11 k-point grid, beyond the $\Gamma$ point which is essential for accurately capturing heat transport properties.
@@ 2.3 Overview of the investigated systems and potentials for benchmarking
The MOFs chosen for benchmarking the phonons calculated with the MACE-MP-MOF0 models are widely studied and hence, enables thorough comparison of performance across the aforementioned models and experiments reported in literature. The investigated systems of MOF-5 (Zn), UiO-66 (Zr), MOF-74 (Zn) and MIL-53 (Al) are diverse in their nodes, linkers and topology (refer Table 2). In this work, we study the large-pore configuration of the flexible framework, MIL-53 (Al). The presence or absence of these MOFs in the training sets of the two versions of MACE-MP-MOF0 models help analyze the transferability of the model and the performance dependence on the choice of training set. In addition to the primary benchmarking of phonon density of states and band structures for these four MOFs, benchmarking of the phonon-derived bulk modulus is done for several other well-known MOFs in literature which are not a part of the curated dataset for demonstrating the transferability of the MACE-MP-MOF0 model.
@@
MOF-5 (Zn2+) No No
UiO-66 (Zr4+) No Yes
MOF-74 (Zn2+) No Yes
MIL-53 (Al3+) Yes Yes
Table 1: The presence of the investigated MOFs for benchmarking the MACE-MP-MOF0 models in their respective training sets indicated by Yes and absence by No
@@ 2.5 Benchmarking predictions of forces, energies and stress
The analysis of element-wise mean absolute errors (MAE) in energies relative to DFT for the MOFs in the QMOF database containing the 24 elements in the curated dataset reveal that the MACE-MP-0b predictions were 80% larger than the MACE-MP-MOF0-v2 MAE predictions of 0.016 eV/atom (see SI). The MACE-MP-MOF0-v2 model shows significant improvement of 50% in the energy MAE for MOFs with metal nodes like Zn, Al, Mg (which constitute a large set in the QMOF database) and in Hf and Zr-based MOFs (which are widely studied due to their importance in defect engineering of MOFs [43]). The energy MAEs for integral elements of organic linkers like C, H and O common to all MOFs are also approximately lowered by 36% with MACE-MP-MOF0-v2. Hence, overall we achieve a significant improvement in the MAE for the majority of the 24 elements in the curated MOFs, which cover 60% of the QMOF database, by including a small number of their reference configurations in the training dataset.
In addition to benchmarking the MACE-MP-MOF0 models for the curated dataset, we test the model for forces and energies on 70 more diverse MOFs unseen by the model during training (see SI for structure details). As shown in Table 4, the MACE-MP-MOF0 models exhibit energy RMSDs relative to DFT that are five times lower than those of MACE-MP-0b, while also delivering 30% more accurate forces and stresses for these out-of-sample MOFs. We observe the deviations on out-of-sample MOFs to be 10 times the RMSDs of MACE-MP-MOF0 models on the curated dataset in Table 3, which demonstrates the significant improvement that can be achieved by adding a few reference configurations of MOFs to the training DFT data.
@@ 2.6 Benchmarking vibrational properties
After benchmarking forces, energies and stresses we quantitatively analyze the derived phonon density of states and band structures. In Figure 4, phonon density of states (DOS) data is shown up to 20THz (see SI for the full range of predicted frequencies). The MACE-MP-MOF0 and MACE-MP-MOF0-v2 models are in very good agreement with DFT DOS, even well above the low frequency range ($\leq$ 6THz). We were able to eliminate all spurious imaginary modes with the MACE-MP-MOF0 models to obtain physically meaningful frequencies. While GFN1-xTB and MACE-MP-0b capture certain aspects of the DOS curves accurately,, they fail to accurately capture the vibrational properties of the most prototypical MOFs like MOF-5 and UiO-66.
@@
The similar errors obtained with MACE-MP-MOF0 and MACE-MP-MOF0-v2 in Table 5 show that the absence or presence of the investigated systems in their respective training sets does not significantly affect the performance of the model, and that the curated dataset is representative to capture the lattice dynamics of out-of-training set MOFs. Since MACE-MP-MOF0 and MACE-MP-MOF0-v2 models exhibit similar frequency errors, we use the slightly better performing MACE-MP-MOF0-v2 model for these MOFs to show the accurate overlap between the band structures predicted with DFT and MACE-MP-MOF0-v2 in Figure 5. Due to the large number of bands in MOFs, the band structures in figure 5 are restricted to the low frequency region.
Figure 4: Density of states (DOS) in the frequency range ($\leq$ 20 THz) for the investigated MOFs obtained with the linear tetrahedron method with a pitch of 0.01 THz predicted by the respective methods overlaid on DFT data.
MOF $RMSD_{\text{MACE-MP-MOF0-v2}}$ $RMSD_{\text{MACE-MP-MOF0}}$ $RMSD_{\text{MTP}}$ $RMSD_{\text{VASP MLP}}$
@@
MOF-5 0.033 0.039 0.099 0.27
UiO-66 0.082 0.122 0.111 0.312
MOF-74 0.173 0.243 0.093 0.126
MIL-53 0.138 0.126 0.156 0.264
Table 5: Root Mean Square Deviations (RMSDs) of the phonon frequencies in the full range for the investigated MOFs with MTP and VASP MLP data as reported in [22]
@@ 2.7 Benchmarking of mechanical properties
Here, we analyze mechanical properties, such as the bulk modulus, for additional widely studied MOFs beyond the four systems investigated, comparing them to experimental and DFT data reported in the literature. The bulk modulus is obtained from fitting the Birch-Murnaghan equation of state to energy-volume data obtained by straining the cell ±$\pm 2\%$. From Figure 6, we observe that MACE-MP-MOF0 and MACE-MP-MOF0-v2 are able to qualitatively capture bulk moduli trends, as compared to DFT and experimental data, as well as quantitatively reproduce the values with minor deviations.
The deviations in bulk modulus with MACE-MP-MOF0 versions relative to DFT are the largest for MIL-125 in Figure 6 b), which notably, is the only Ti-based MOF in the curated dataset. Serendipitously, the MACE-MP-MOF0 and MACE-MP-MOF0-v2 predictions of bulk moduli actually perform better than the explicit DFT data, relative to experiments. The DFT data includes VdW corrections, which tends to overestimate the long range interactions in MOFs [46] and therefore produces higher bulk modulus than experiments. As the MACE-MP-MOF0 model was trained on DFT data with vdW corrections included, these long-range interactions are therefore slightly underestimated in the model relative to DFT, which moderates the bulk moduli predictions. Hence, the small RMSDs obtained for phonon frequencies with MACE-MP-MOF0-v2 and MACE-MP-MOF0 translate to low deviations in phonon-derived properties, such as the bulk modulus.
The agreement between experiments, DFT and MACE-MP-MOF0 models in obtaining the bulk modulus of out-of-sample, well-known MOFs like ZIF-4, ZIF-8, IRMOF-14 and UiO-66-Hf demonstrates the transferability of the model.
Figure 6: Bulk modulus MACE-MP-MOF0 and MACE-MP-MOF0-v2 predictions as compared to a) only DFT and b) DFT and Experiments where ZIF-4, ZIF-8, IRMOF-14 and UiO-66-Hf/Ce are out-of-curated dataset MOFS. DFT data is self-generated or reported in [47, 48, 24, 49] and experiment reported in [50, 51, 52, 53, 54]
@@ 2.8 Negative Thermal Expansion in MOFs
The data available in literature for negative thermal expansion (NTE) obtained through experiments for MOFs is scarce. At the same time, obtaining NTE with DFT is challenging due to the expensive quasi-harmonic phonon calculations for MOFs. Hence, in this work, we restrict our analysis of NTE to MOF-5 and UiO-66 as they are widely studied in experiments and their high symmetry and isotropic nature reduces the cost of such expensive computations.
@@ 3 Discussion
Benchmarking the vibrational properties of the investigated MOFs highlights the limitations of MACE-MP-0b and GFN-1xTB, while demonstrating the excellent agreement of the MACE-MP-MOF0 models with DFT in accurately capturing the full phonon vibrational spectra. We further rationalize the performance of these models by analyzing their ability in accurately capturing the atomic positions and interactions in the equilibrium structure 7 for MOF-74. An accurate capture of atomic positions in the MOF provides a better description of forces, which ultimately correlates to the partial phonon density of states (PDOS) obtained with DFT in Figure 7.
Figure 4 for MOF-74 shows that the GFN1-xTB method is in good agreement with DFT for frequencies lower than 6 THz, above which the errors signficantly increase. The overlap is better in the lower frequency range because this region is dominated by the heavy Zn atom vibrations which have low RMSDs in Table 7. The GFN1-xTB and MACE-MP-0b errors in the atomic positions for C and O, dominating the vibrations beyond 6THz, are ten to hundred times larger than the MACE-MP-MOF0 errors. The greater flexibility and higher rotational degrees of freedom of these atoms in the organic linkers are more difficult to capture than the surrounding metal nodes [16].Therefore, the GFN1-xTB method accurately captures the vibrations of the metal nodes, but it inadequately represents the interactions between atoms in the organic linkers for phonon calculations. As the MACE-MP-MOF0 models consistently achieves ten to thousand times lower errors for all the elements present in the MOFs than GFN1-xTB and MACE-MP-0b, we obtain the high agreement between DFT and MACE-MP-MOF0 model predictions throughout the phonon spectra (see SI for the analysis of other MOFs). MACE-MP-MOF0 demonstrates a significant improvement in capturing the covalent interactions in MOFs, addressing the limitations of MACE-MP-0b and GFN1-xTB, which primarily capture non-covalent interactions. This enhancement positions MACE-MP-MOF0 as a promising tool for systems beyond MOFs, where covalent interactions dominate, such as in covalent organic frameworks.[60].
@@
In summary, in this work, we present a fine-tuned MLP for MOFs, MACE-MP-MOF0, that can be used to produce ab-initio quality phonons in a high-throughput way. The model was trained on a representative dataset of 127 MOFs selected by efficiently sampling the phase space based on MACE descriptors which significantly reduced the computational efforts in generating DFT reference data. The MACE-MP-MOF0 model predicts phonon-derived bulk modulus in excellent agreement with experiments and DFT on MOFs unseen by the model, demonstrating its transferability.
@@
The MACE-MP-MOF0 model presented here is 50% faster than the foundation model (with dispersion corrections included) and 10 times more accurate in obtaining optimized geometries, forces, energies and stresses which are crucial for obtaining accurate vibrational properties. The MACE-MP-MOF0 model is 90% faster than GFN1-xTB when handling large MOFs (500 atoms per unit cell) (see SI) The computational efficiency and accuracy of the model make it an excellent candidate for screening phonon-derived properties, such as bulk modulus, in MOFs, providing a quantitative description that aligns closely with experimental and DFT results. The model can calculate phonons within the quasi-harmonic approximation to obtain ab initio-level thermal expansions, which would otherwise be computationally expensive to obtain using DFT for MOFs.
@@
Through the preliminary analysis for thermal expansion of MOF-5 and UiO-66 with MACE-MP-MOF0 and DFT, we conclude that the quasi-harmonic approximation does not provide a sufficient description of anharmonic effects in MOFs and is an area with scope for improvement in the computational study of vibrational properties of MOFs. While MACE-MP-MOF0 accurately predicts phonons in the majority of the frequency range (0 to 60 THz) which dominate most of the physical properties of interest, nevertheless, further improvement is possible by capturing the high frequency range (90 - 100 THz, which correspond to the vibration of H atoms in the linkers) as shown in Figure 8. We would also like to highlight that while the current model covers 60% of MOFs in the QMOF database that span the same chemical space of metal nodes as the 127 MOFs, these sampled MOFs are majority closed-shell metal ions, which avoids electronic spin-degrees of freedom. In MOFs with magnetic elements, having training data of atomic configurations with different spin-states in the same MOF can lead to poor training of the MLP. HKUST-1, a well-known Cu-based MOF, is the only MOF in the curated dataset with a magnetic element which contributes to the higher element-wise MAE obtained with MACE-MP-MOF0 for Cu (refer SI). In addition to the model presented here being ready-to-use for a wide range of metallic nodes, organic linkers, and topologies, it is significantly easier to re-parameterize than traditional force fields like MOF-FF for adding new species by including a few reference configurations of the MOF to the DFT training set.

### v3

@@ 1 Introduction
Semi-empirical Quantum Mechanical methods, such as the Density Functional Tight Binding (DFTB) method, have been used to compute phonons in few highly symmetric isoreticular MOFs [14]. Several DFTB schemes have been developed, with DFTB3 [15] being the most widely used. While DFTB is 2–3 orders of magnitude faster than DFT, its pair- wise parametrization strategy and lack of available metal atom parameters limits its use in screening applications, particularly pertaining to phonons in MOFs. A notable variant of DFTB3, GFN1-xTB [16] follows a element-specific parameters strategy and extends its applicability to nearly the entire periodic table, up to thousands of atoms. In [17], the GFN1-xTB method has been reported to produce cell parameters within a 5% deviation (relative to experiments) for 75% of the MOFs in the CoRE database,[18, 19] showing promise for adsorption applications and obtaining binding energies. However, GFN1-xTB still scales quadratically with the number of electrons which may present challenges for larger MOFS, and has not been used for the analysis of vibrational properties in MOFs.
Traditional force fields like UFF [20] and CHARMM [21] are widely used due to their scalability and significant speed up in calculations compared to tight-binding DFT. UFF4MOF [22], the UFF parametrization on MOFs, can be used for rapid structure prediction and screening. However, this transferability comes with limitations in accurately predicting dynamical properties such as phonons. [23, 22]. Other force fields, like the MOF-FF model [24] also predicted accurately the lattice parameters of some well known MOFs like MOF-5, UiO-66 within 0.5% deviation relative to DFT. Force fields derived with a focus on vibrational properties for MOFs have also been reported like VMOF [25]. While the VMOF model reproduced optimized lattice parameters and phonon density of states in good agreement with DFT, important phonon-derived properties like bulk modulus obtained with VMOF underestimated DFT predictions by more than 50% even for standard MOFs like UiO-66 and MOF-5 [25]. These limitations of such classical force fields are because of the vast combinations of possible frameworks in the MOF chemical space that make their parameter optimization and selection of functional form extremely challenging.
@@ 2.1 Dataset Curation
For sampling these 108 structures, we considered predominantly non-spin polarized MOFs with pore limiting diameters (PLDs) greater than 3.6 Å, using nitrogen as a probe molecule. MACE [35] descriptors hold a large amount of information regarding the atomic features of MOFs and hence were used to sample 100 diverse MOFs. These MOFs were selected to span over all the 7 crystal symmetry systems and diverse bonding interactions. Therefore, as shown in Figure 1 (c, d), this dataset consists of 127 representative MOFs spread over a wide range of 24 elements in the inorganic clusters and organic ligands. Figure 1 (a, b) shows the MACE descriptor space diversity for the curated 127 MOFs (refer the SI for the main features of the sampled MOFs).
@@ 2.2 Phonon Workflow
The Harmonic Approximation (HA) for computing phonons fails to capture essential physical properties for MOFs such as thermal expansion, phase transition and elastic moduli [43]. On the other hand, the Quasi-Harmonic approximation (QHA) is an extension of the HA that aims to capture anharmonicity by introducing the volume dependence of frequencies due to temperature and pressure effects. In this work, we use the Phonopy package [44] version 2.29.0 to study thermal and mechanical properties of MOFs by studying the lattice dynamics under the QHA. A $11\times 11\times 1$ 11 k-point grid, beyond the $\Gamma$ point, is consistently used for all phonon calculations, which is essential to accurately capture heat transport properties.
@@ 2.3 Overview of the investigated systems and potentials for benchmarking
The MOFs chosen for benchmarking the phonons calculated with the MACE-MP-MOF0 models are widely studied and hence, enables thorough comparison of performance across the aforementioned models and experiments reported in literature. The investigated systems of MOF-5 (Zn), UiO-66 (Zr), MOF-74 (Zn) and MIL-53 (Al) are diverse in their nodes, linkers and topology (refer Table 2). In this work, we study the large-pore configuration of the flexible framework, MIL-53 (Al). The number of reference configurations of these MOFs in the training sets of the two versions of MACE-MP-MOF0 models help analyze the transferability of the model and the performance dependence on the choice and size of training set. In addition to the primary benchmarking of phonon density of states and band structures for these four MOFs, benchmarking of the phonon-derived bulk modulus is done for several other well-known MOFs in literature which are not a part of the curated dataset for demonstrating the transferability of the MACE-MP-MOF0 model.
@@
MOF-5 (Zn2+) 114 92
UiO-66 (Zr4+) 68 58
MOF-74 (Zn2+) 59 60
MIL-53 (Al3+) 30 29
Table 1: The number of reference structures generated for the investigated MOFs for benchmarking the MACE-MP-MOF0 models in their respective training sets
@@ 2.5 Benchmarking predictions of forces, energies and stress
The analysis of element-wise mean absolute errors (MAE) in energies relative to DFT for the MOFs in the QMOF database containing the 24 elements in the curated dataset reveal that the MACE-MP-0b predictions were 80% larger than the MACE-MP-MOF0-v2 MAE predictions of 0.016 eV/atom (see SI). The MACE-MP-MOF0-v2 model shows significant improvement of 50% in the energy MAE for MOFs with metal nodes like Zn, Al, Mg (which constitute a large set in the QMOF database) and in Hf and Zr-based MOFs (which are widely studied due to their importance in defect engineering of MOFs [45]). In this subset of the QMOF database, more than 75% of MOFs show improved predictions, with the remaining 25% primarily coming from the Cu-MOFs. The energy MAEs for integral elements of organic linkers like C, H and O common to all MOFs are also approximately lowered by 36% with MACE-MP-MOF0-v2. Hence, overall we achieve a significant improvement in the MAE for the majority of the 24 elements in the curated MOFs, which cover 60% of the QMOF database, by only including a small number of their reference configurations in the training dataset.
In addition to benchmarking the MACE-MP-MOF0 models for the curated dataset, we evaluate its performance on forces, as well as energies, on a set of 70 more diverse MOFs unseen by the model during training (see SI for structure details). As shown in Table 4, the MACE-MP-MOF0 models exhibit energy RMSDs relative to DFT that are five times lower than those of MACE-MP-0b, while also delivering 30% more accurate forces and stresses for these out-of-sample MOFs. We observe the deviations on out-of-sample MOFs to be 10 times the RMSDs of MACE-MP-MOF0 models on the curated dataset in Table 3, which demonstrates the significant improvement that can be achieved by adding a few reference configurations of MOFs to the training DFT data.
@@ 2.6 Benchmarking vibrational properties
After benchmarking forces, energies and stresses we quantitatively analyze the derived phonon density of states and band structures. In Figure 4, phonon density of states (DOS) data is shown up to 20THz (see SI for the full range of predicted frequencies). The MACE-MP-MOF0 and MACE-MP-MOF0-v2 models are in very good agreement with DFT DOS, even well above the low frequency range ($\leq$ 6THz). We were able to eliminate all spurious imaginary modes with the MACE-MP-MOF0 models to obtain physically meaningful frequencies. While GFN1-xTB and MACE-MP-0b capture certain aspects of the DOS curves accurately, they fail to accurately capture the vibrational properties of the most prototypical MOFs like MOF-5 and UiO-66.
@@
The similar errors obtained with MACE-MP-MOF0 and MACE-MP-MOF0-v2 in Table 5 show that the different reference configurations of the investigated systems in their respective training sets does not significantly affect the performance of the models. Since MACE-MP-MOF0 and MACE-MP-MOF0-v2 models exhibit similar frequency errors, we use the slightly better performing MACE-MP-MOF0-v2 model for these MOFs to show the accurate overlap between the band structures predicted with DFT and MACE-MP-MOF0-v2 in Figure 5. Due to the large number of bands in MOFs, the band structures in figure 5 are restricted to the low frequency region.
Figure 4: Density of states (DOS) in the frequency range ($\leq$ 20 THz) for the investigated MOFs obtained with the linear tetrahedron method with a pitch of 0.01 THz predicted by the respective methods overlaid on DFT data. DFT DOS were calculated from force constants available in [49, 23] on a $11\times 11\times 1$ 11 mesh
MOF $Full\_Mesh_{\text{MACE-MP-MOF0-v2}}$ $Full\_Mesh_{\text{MACE-MP-MOF0}}$ $\Gamma-only_{\text{MTP}}$ $\Gamma-only_{\text{VASP MLP}}$
@@
MOF-5 0.033 (0.085) 0.039 (0.083) 0.099 0.27
UiO-66 0.082 (0.186) 0.122 (0.198) 0.111 0.312
MOF-74 0.173 (0.427) 0.243 (0.436) 0.093 0.126
MIL-53 0.138 (0.257) 0.126 (0.239) 0.156 0.264
Table 5: Comparison of Root Mean Square Deviations (RMSDs) of the phonon frequencies in the full range for the investigated MOFs with MTP and VASP MLP data as reported in [23]. $\Gamma$-only errors with MACE-MP-MOF0 models are provided in brackets.
@@ 2.7 Benchmarking of mechanical properties
The bulk modulus serves as a sensitive and computationally efficient metric for assessing phonon accuracy, as both phonon frequencies and bulk modulus arise from second derivatives of energy- phonons with respect to atomic displacements and bulk modulus with respect to volume. Since both properties share similar force accuracy requirements, an accurate bulk modulus indicates that the elastic response and interatomic potential curvature are well captured, suggesting reasonable phonon predictions. Hence, in this section, we analyze mechanical properties, such as the bulk modulus, for additional widely studied MOFs beyond the four systems investigated, comparing them to experimental and DFT data reported in the literature. The bulk modulus is obtained from fitting the Birch-Murnaghan equation of state to energy-volume data obtained by straining the cell ±$\pm 2\%$. From Figure 6, we observe that MACE-MP-MOF0 and MACE-MP-MOF0-v2 are able to qualitatively capture bulk moduli trends, compared to DFT and experimental data, as well as quantitatively reproduce the values with minor deviations.
The deviations in bulk modulus with MACE-MP-MOF0 versions relative to DFT are the largest for MIL-125 in Figure 6 b), which notably, is the only Ti-based MOF in the curated dataset. We hypothesize that the deviation is due to insufficient training data for Ti-based MOFs. Serendipitously, the MACE-MP-MOF0 and MACE-MP-MOF0-v2 predictions of bulk moduli actually perform better than the explicit DFT data, relative to experiments. The DFT data includes VdW corrections, which tends to overestimate the long range interactions in MOFs and therefore produces higher bulk modulus than experiments [50]. As the MACE-MP-MOF0 model was trained on DFT data with vdW corrections included, these long-range interactions are therefore slightly underestimated in the model relative to DFT, which moderates the bulk moduli predictions. Hence, the small RMSDs obtained for phonon frequencies with MACE-MP-MOF0-v2 and MACE-MP-MOF0 translate to low deviations in phonon-derived properties such as the bulk modulus.
The agreement between experiments, DFT and MACE-MP-MOF0 models in obtaining the bulk modulus of several diverse out-of-sample, well-known MOFs like ZIF-4, ZIF-8, UiO-66-Ce, UiO-66-Hf and promising MOFs for direct air capture like XEDPON, SUSZOW and XEXMEU [51], thus demonstrates the transferability and accuracy of the model in obtaining bulk modulus of MOFs unseen by the model during training.
Figure 6: Bulk modulus predictions of MACE-MP-MOF0 models as compared to a) only DFT and b) DFT and Experiments where ZIF-4, ZIF-8, XEDPON, XEXMEU, SUSZOW [51], UiO-66-Hf/Ce are diverse out-of-curated dataset MOFS. DFT data is self-generated or reported in [52, 53, 25, 54] and experiment reported in [55, 56, 57, 58, 59]
@@ 2.8 Negative Thermal Expansion in MOFs
Finally, we evaluated our model on negative thermal expansion (NTE). Experimental data on NTE for MOFs is limited, making it challenging to benchmark against a wide range of models and structure. In addition, obtaining NTE with DFT is computationally expensive due to the quasi-harmonic phonon calculations required for MOFs. To address this challenge, in this work, we focused our NTE analysis on MOF-5 and UiO-66 as they are widely studied in experiments and their high symmetry and isotropic nature reduces the cost of such expensive computations.
@@ 3 Discussion
Benchmarking the vibrational properties of the investigated MOFs highlights the limitations of MACE-MP-0b and GFN-1xTB, while demonstrating the excellent agreement of the MACE-MP-MOF0 models with DFT in accurately capturing the full phonon vibrational spectra. We further rationalize the performance of these models by analyzing their ability in accurately capturing the atomic positions and interactions in the equilibrium structure in Table 7 for MOF-74. An accurate capture of atomic positions in the MOF provides a better description of forces, which ultimately correlates to the partial phonon density of states (PDOS) obtained with DFT in Figure 7.
Figure 4 for MOF-74 shows that the GFN1-xTB method is in good agreement with DFT for frequencies lower than 6 THz, above which the errors significantly increase. The overlap is better in the lower frequency range because this region is dominated by the heavy Zn atom vibrations which have low RMSDs in Table 7. The GFN1-xTB and MACE-MP-0b errors in the atomic positions for C and O, dominating the vibrations beyond 6THz, are ten to hundred times larger than the MACE-MP-MOF0 errors. The greater flexibility and higher rotational degrees of freedom of these atoms in the organic linkers are more difficult to capture than the surrounding metal nodes [17].Therefore, the GFN1-xTB method accurately captures the vibrations of the metal nodes, but it inadequately represents the interactions between atoms in the organic linkers for phonon calculations. As the MACE-MP-MOF0 models consistently achieves ten to thousand times lower errors for all the elements present in the MOFs than GFN1-xTB and MACE-MP-0b, we obtain the high agreement between DFT and MACE-MP-MOF0 model predictions throughout the phonon spectra (see SI for the analysis of other MOFs). MACE-MP-MOF0 demonstrates a significant improvement in capturing the covalent interactions in MOFs, addressing the limitations of MACE-MP-0b and GFN1-xTB, which primarily capture non-covalent interactions. This enhancement positions MACE-MP-MOF0 as a promising tool for systems beyond MOFs, where covalent interactions dominate, such as in covalent organic frameworks.[65].
@@
In summary, in this work, we present a fine-tuned MLP for MOFs, MACE-MP-MOF0, that can be used to produce ab initio quality phonons in a high-throughput way. The model was trained on a representative dataset of 127 MOFs selected by efficiently sampling the phase space based on MACE descriptors which significantly reduced the computational efforts in generating DFT reference data. The MACE-MP-MOF0 model predicts phonon-derived bulk modulus in excellent agreement with experiments and DFT on MOFs unseen by the model, demonstrating its transferability.
@@
The MACE-MP-MOF0 model presented here is also 50% faster than the MACE-MP-0b foundation model (with dispersion corrections included) and 10 times more accurate in obtaining optimized geometries, forces, energies and stresses which are crucial for obtaining accurate vibrational properties. Additionally, MACE-MP-MOF0 model is 90% faster than GFN1-xTB for large MOFs (500 atoms per unit cell) (see SI) The computational efficiency and accuracy of the model make it an excellent candidate for screening phonon-derived properties, such as bulk modulus, in MOFs, providing a quantitative description that aligns closely with experimental and DFT results. The model can calculate phonons within the quasi-harmonic approximation to obtain ab initio-level thermal expansions, which would otherwise be computationally expensive to obtain using DFT for MOFs.
@@
Through the preliminary analysis for thermal expansion of MOF-5 and UiO-66 with MACE-MP-MOF0 and DFT, we conclude that the quasi-harmonic approximation does not provide a sufficient description of anharmonic effects in MOFs and is an area with scope for improvement in the computational study of vibrational properties of MOFs. While MACE-MP-MOF0 accurately predicts phonons in the majority of the frequency range (0 to 60 THz) which dominate most of the physical properties of interest, nevertheless, further improvement is possible by capturing the high frequency range (90 - 100 THz, which correspond to the vibration of H atoms in the linkers) as shown in Figure 8. We would also like to highlight that while the current model covers 60% of MOFs in the QMOF database that span the same chemical space of metal nodes as the 127 MOFs, these sampled MOFs are majority closed-shell metal ions, which avoids electronic spin-degrees of freedom. In MOFs with magnetic elements, having training data of atomic configurations with different spin-states in the same MOF can lead to poor training of the MLP. HKUST-1, a well-known Cu-based MOF, is the only MOF in the curated dataset with a magnetic element which contributes to the higher element-wise MAE obtained with MACE-MP-MOF0 for Cu (refer SI). In addition to the model presented here being ready-to-use for a wide range of metallic nodes, organic linkers, and topologies, it is also significantly easier to re-parameterize for new species by simply including a few reference configurations of the MOF into the DFT training set, compared to traditional force fields.